An electric heater with a power rating of 1,50 kW heated an empty greenhouse containing only moist air.

The volume of the room is 684.00m^3, the density (d) of dry air is 1.225kg/m^3.

The specific heat capacity of dry air is S(air) = 1.012J g ^-1 degrees ^-1

What is the change in temperature of the greenhouse after heating for 230 minutes?
1
Expert's answer
2020-12-07T03:16:59-0500

Power, P = 1.5kW = 1500W

Volume, v = 684m³

Density, d = 1.225kg/m³

Time, t = 230mins = 13800s

SHC, c of dry air = 1.012J/g°c


∆T = Q/m*c


Q= Pt = 1500*13800 = 20700000J

m = dv = 1.225*684 = 837.9kg = 837900g


∴ ∆T = 20700000 / (837900*1.012)

= 20700000/847954.8

= 24.4°C